if 10 ml of NaOH requires 8 ml of 2M of HCl to neutralise. then 10ml of NaOH will require how much quantity of 2M of H2SO4

Grade:10

1 Answers

Vikas TU
7 years ago
M1V1*v.f = M2V2*v.f
2*8*1 = 10*M*1
M = 16/10 = 1.6
 
Now,
1.6*10*1 = 2*V*2 {Valence factor for H2SO4 is 2}
V = 4 ml.